OpenStudy (anonymous):

How far above the surface of Earth do you need to go for g to be reduced to 0.5 of its value on the surface of Earth?

OpenStudy (anonymous):

g (at surface)=GM/R^2 (where, R= Radius of earth- 6400 km, M =mass of earth) g (at height r) =GM/r^2 But GM/r^2=1/2*GM/R^2 Or r=R*sqrt(2) =6400*sqrt(2)